I would like some help in determining all of the real and imaginary zeros of x^4+3x^2-4.
Thank you for your help

Answer by tutorcecilia(2152)

x^4+3x^2-4 [factor]
(x^2-1)(x^2+4) [re-factor the difference of two squares]
(x+1)(x-1)(x^2+4) [set each factor equal to zero]
x+1=0
x=-1
.
x-1=0
x=1
.
and
x^2+4=0
x^2=-4
sqrt(x^2)=sqrt(4)
x=4i (an imaginary number]
.
So, x=-1,1,-4i, or +4i